OpenStudy (anonymous):

jan can mow her neighbor's lawn in 3hrs. Ian can mow the sam lawn in 2hrs. to the nearest minute,how long will it take if they mow together????

OpenStudy (anonymous):

guess what? \[\frac{3\times 2}{3+2}\]

OpenStudy (anonymous):

its the same as the other?!

OpenStudy (anonymous):

same as all of them adam can do the job in A hours, eve can do it in E together they can do it in \[\frac{A\times E}{A+E}\] hours